Grade 11Physical Chemistry

60 g of a compound on analysis gave C = 24 g H=4 g ) = 32 g Its empirical formula is

akash
8 Years agoGrade 11

1 Answer

poonam
8 Years ago
Assuming that the third element is oxygen
carbon=24/12=2
hydrogen=4/1=4
oxygen=32/16=2
so the empirical formula is c2h4o2